# Policies for multiplexed quantum repeaters: theory and practical performance analysis

### v1

@@ VIII Conclusions and directions for future work
An important consideration for practical and scalable implementation of entanglement distribution policies is the classical communication (CC) cost associated with non-deterministic entanglement swapping. Here, we have considered these overheads and shown that not only do multiplexed policies retain their advantage over non-multiplexed policies when CC costs are included, but this advantage can also increase with an increasing number of nodes. This is a surpising and counter-intuitive result, and shows that policies that use some knowledge of the network state (but not necessarily full, global knowledge) can enhance network performance even when CC costs associated with such knowledge are accounted for. This is a very important conclusion from the point of view determining useful policies beyond the fully-local ones, especially for large quantum networks. Our work provides much needed evidence in support of the widely believed conjecture that policies that utilize knowledge of the network state can outperform policies that do not use such information.
@@
Finally, to make the numerical and analytical analyses of this work more concrete, we performed an experimental demonstration of multiplexing using a high-dimensional biphoton frequency comb (BFC), which would form the backbone of our proposed experimental implementation of a linear-chain quantum network with multiplexing capabilities. We then assessed the anticipated performance of such a network over 100 km for two concrete memory platforms, namely rare-earth ion and diamond vacancy based quantum memories, when using our multiplexing-based policies.
A universal principle that is evident from the results of this work is that policies must be tailored to the hardware parameters of the network. This motivates us to further explore the question of policies, given that no single entanglement distribution policy can provide an advantage in all network scenarios.
@@ Acknowledgements.
The authors thank Thomas Searles and Sanjaya Lohani for helpful discussions. This work was supported by the Army Research Office Multidisciplinary University Research Initiative (ARO MURI) through the grant number W911NF2120214. SH also acknowledges support from the RCS program of Louisiana Boards of Regents through the grant LEQSF(2023-25)-RD-A-04, NSF grant PHY-2110273, funds from Xairos Inc., and funds from the Hearne Institute for Theoretical Physics. PB and HL also acknowledge the support of the U.S. Air Force Office of Scientific Research as well as the US-Israel Binational Science Foundation. SK acknowledges financial support from the German BMBF (Hybrid).
Appendix A Noise model and decoherence
@@
For a proof, we refer to Ref. [43, Appendix E]. This implies that the fidelity of the state after $m$ time steps is $\frac{1}{4}(1+\mathrm{e}^{-2m/m_{1}^{\star}}+2\mathrm{e}^{-2m/m_{2}^{\star}})$.
@@
In other words, we consider equal dephasing and amplitude damping. The fidelity after $m$ time steps is then equal to
$f(m)=\frac{1}{4}(1+3\mathrm{e}^{-\frac{m}{m^{\star}}}).$ (23)
Note that we have chosen the value of $m_{2}^{\star}$ in Eq. (22) such that, at the cutoff time $m^{\star}$, the fidelity of the entangled qubit pair is $f(m^{\star})=\frac{1}{4}(1+3/e)\approx 0.5259$. We emphasize that our results can be applied to other choices of $m_{1}^{\star}$ and $m_{2}^{\star}$—in particular, choices that could take dephasing as the dominant source of noise [94] or in order to make connections to prior works [58, 20, 23, 41, 30].
@@
$(a+b)\Phi^{+}+(a-b)\Phi^{-}+c\Psi^{+}+c\Psi^{-}.$ (25)
As such, its entanglement can be characterized entirely by the quantity $a+b$, namely, the fidelity of the state with respect to $\Phi^{+}$. In particular, the state is entangled if and only if this fidelity is strictly greater than $\frac{1}{2}$; see, e.g., [28, Chapter 2]. Therefore, under the assumptions in Eq. (22), such that the fidelity is given by Eq. (23), we find that the state in Eq. (21) is entangled for all time steps $m\in\{0,1,2,\dotsc,m^{\star}\}$. Furthermore, inserting $F=\frac{1}{2}$ in Eq. (24), we find that the age at which entanglement is lost is $\lceil m^{\star}\log(3)\rceil$.
Appendix B Entanglement distillation with the BBPSSW protocol
@@
With this expression for the fidelity after entanglement distillation, if we let $f_{1}=f(m_{1})$ and $f_{2}=f(m_{2})$, with $m_{1}$ and $m_{2}$ the ages of the two entangled states, then we can calculate the updated age as $m^{\prime}=f^{-1}(F_{\text{distill}}(f(m_{1}),f(m_{2})))$, with the expression for $f^{-1}$ given in Eq. (24). In particular, the updated age $m^{\prime}$ is given by:
@@
Now, distillation should be considered useful only when it can improve upon the fidelity of the best of the two links. In other words, in order for distillation to be useful, we should have $F_{\text{distill}}(f_{1},f_{2})>\max\{f_{1},f_{2}\}$. In Fig. 17, we plot the $(f_{1},f_{2})$ region defined by this inequality.
B.1 The pumping policy
@@
B.2 Banding vs. pumping
@@
Figure 18: Comparison of the swap-distill (top) and distill-swap (bottom) policies for a three-node chain with $n_{ch}$ links connecting the neighboring nodes, as described in Sec. B.2. Here we use the banded policy to perform distillation. The swap-distill policy is not useful beyond initial age $m^{\star}/2$, whereas distill-swap can extend this threshold to $m^{\star}$.
@@
Figure 19: Comparison of the swap-distill (top) and distill-swap (bottom) policies for a three-node chain with $n_{ch}$ links connecting the neighboring nodes, as described in Sec. B.2. Now, instead of the banding policy, we use the pumping policy for distillation. As with banded purification, swap-distill is not useful beyond an initial age of $m^{\star}/2$, whereas distill-swap can extend this threshold, but now not quite up to $\approx m^{\star}$. This is a consequence of the fact that the pumping policy has a strict limit on the fidelity that can be achieved with an increasing number of channels, as shown in Eq. (5).
@@
Appendix C Distill-swap vs swap-distill policy: analytical results
Here we analytically compare the performance of the distill-swap and swap-distill policies in the case of three nodes. To simplify the analysis, we restrict ourselves to the cases with two and four channels, and consider all elementary links to be active simultaneously with the same fidelity. The general result for an arbitrary number of channels is difficult to calculate analytically given the large number of possible trajectories that can lead to the successful generation of at least one end-to-end link. Seeing in Appendix B.2 that the banded policy performs better than pumping, in this section we look exclusively at the banded policy. The performance of the policies is evaluated by the success probability of establishing at least one end-to-end entangled link and by the average fidelity of an end-to-end link.
C.1 Two-channel case
Let $f_{0}$ be the initial fidelity of both the active links on each side given in terms of the initial age $m_{0}$ according to Eq. (23) as $f_{0}=f(m_{0})=\frac{1}{4}(1+3\mathrm{e}^{-m_{0}/m^{\star}})$. Let us first consider the distill-swap policy. The probability of distillation being successful simultaneously on both sides followed by a successful entanglement swapping of the distilled links is given by
@@ C.1 Two-channel case
C.2 Four-channel case
@@
Figure 20: Comparison of the distill-swap and swap-distill policies for a three-node chain with four channels connecting the neighboring nodes, as described in Sec. C. We look at the probability of successfully producing at least one end-to-end entangled link (top) and the average age of that link (when successfully produced) (bottom). For distillation, we use the banded policy. We see again that the swap-distill policy is not useful beyond and initial age of $m^{\star}/2$, whereas distill-swap can extend this threshold up to $m^{\star}$ by using an increasing number of channels.
In Fig. 20, we show the results for four channels. It is clear from Fig. 20 that swap-distill has a higher probability of success compared to distill-swap when the initial age of the links is less than $m^{\star}/2$, beyond which it fails to succeed. Unlike distillation, success probability of entanglement swapping is independent of the quality of the used links. This is the primary reason for an asymmetry in the success rates of swap-distill and distill-swap policies. The probability of at least one entanglement swap being successful increases exponentially with the number of attempts available, whereas the probability of all distillations being successful falls with an increasing number of attempts, irrespective of the ages of the links involved. Thus, it is clear that if the aim is to generate at least one end-to-end link as quickly as possible, i.e., if the aim is to minimize waiting time, the swap-distill is the policy of choice. At the same time, distill-swap succeeds until a larger initial age threshold; thus, if the elementary links are of low fidelity to begin with, distilling first might be judicious. Also, in terms of fidelity enhancement (age reduction), distill-swap performs better. These conclusions are also consistent with the results in Fig. 18 and 19.
Appendix D Other policies for multiplexing based on swap-asap

### v2

@@ VII Conclusions and directions for future work
We have shown that not only do our quasi-local multiplexed policies retain their advantage over fully-local policies when CC costs are included, but this advantage can also increase with an increasing number of nodes. This is a surprising and counter-intuitive result, and shows that policies that use some knowledge of the network state (but not necessarily full, global knowledge) can enhance network performance, even when the CC costs associated with such knowledge are accounted for. The advantage attained by our quasi-local policies is rooted in the fact that they only require CC between connected regions of the chain; therefore, for most time steps, end-to-end CC is not required. This is an important conclusion from the point of view determining useful policies beyond the fully-local ones, especially for large quantum networks. We also benchmarked our policies against the widely studied doubling policy. We have shown via simulations that the FN and SN swap-asap policies can yield a considerable advantage over the doubling policy, both in terms of reducing the average end-to-end waiting time and increasing the fidelity of the end-to-end link. These advantages occur in the most relevant parameter regimes for near-term quantum networks, which correspond to the most resource-constrained settings.
We then considered policies with entanglement distillation, and we proposed a new policy that we call distill-asap. This policy combines the benefits of existing distillation policies, like the banded, greedy, and pumping approaches, and thus is able to outperform the doubling policy nested with distillation. We also provide answers to two important policy questions related to distillation: When is distillation useful, and when it is useful, should we distill then swap, or the other way around? In this direction, the next question of immediate interest for future work is: How much to distill? Indeed, in this work, we only considered entanglement distillation policies that take $N$ links and distill them to one, but one could instead distill $N$ links to some $K$ links greater than one, and such policies have been the study of recent work [99].
@@
Finally, we performed an experimental demonstration of multiplexing using a high-dimensional biphoton frequency comb (BFC), which would form the backbone of our proposed experimental implementation of a linear-chain quantum network with multiplexing capabilities. We then assessed the anticipated performance of such a network over 100 km for two concrete memory platforms, namely rare-earth ion and diamond vacancy based quantum memories, when using our multiplexing-based policies.
Moving forward, it would be interesting to assess the optimality of the policies presented in this work. Optimal policies for multiplexed quantum repeater chains can be obtained using reinforcement learning (RL), using the methods developed in Refs. [44, 43, 45]. By adding the appropriate classical communication costs to such policies, we could assess not only the optimality of the FN and SN swap-asap policies presented in this work, but also we could determine whether quasi-local policies in general can outperform fully-global ones, when CC costs are accounted for.
Data availability
The data generated through the simulations performed in this work and the experimental data that support the findings of this study are available on request from the corresponding author.
Code availability
The custom code developed to perform the simulations in this work that support the findings of this study are available on request from the corresponding author.
Author contributions
SH, SK, HL, and PB conceived the main idea of the work. SK, PB, and SH developed the theory and performed the computations. SK and SH performed the analytical calculations. XC and KC performed the experiments supervised by CW. BK encouraged SH to investigate the role of classical communication. CW, HL, BK and SK supervised the findings of this work. All authors discussed the results and contributed to the writing of the manuscript.
Competing interests
The authors declare no competing interests.
@@ Acknowledgements.
The authors thank Thomas Searles and Sanjaya Lohani for helpful discussions. This work was supported by the Army Research Office Multidisciplinary University Research Initiative (ARO MURI) through the grant number W911NF2120214. SH also acknowledges support from the RCS program of Louisiana Boards of Regents through the grant LEQSF(2023-25)-RD-A-04. PB and HL also acknowledge the support the US-Israel Binational Science Foundation. SK acknowledges financial support from the German BMBF (Hybrid).
Appendix A Properties of our multiplexing policies
In this section, we evaluate our multiplexing policies with respect to the end-to-end waiting time and age of the youngest link. We exclude classical communication costs here.
In Fig. 13, and for all of the plots shown in this paper, the parameter values that we choose are based on what is computationally feasible, but they are also based on what we expect with actual hardware, as we outline in Sec. VI.1. The choice of network parameters such as $p_{\ell},m^{\star}$ etc. in specific examples is also motivated by the intent to show the most prominent and overarching trends.
Of course, FN, SN, and random are not the only three possible options for multiplexing-based policies. For example, one could develop a ranking scheme based on a combination of age and distance, or one could perform a search over all possible pairings of links, looking for the optimal pairings in terms of some figure of merit (such as the sum of lengths of all virtual links formed). With the respect to such optimal ranking schemes, it is not clear how close the SN and FN schemes are to being optimal; they are, however, simple and intuitive policy choices that require no optimization in terms of pairing of links for entanglement swapping, and hence are practical for near-term implementation. We defer the discussion of other policy alternatives and comments on the optimality of SN and FN swap-asap to Appendix E.
Let us now compare our policies with the doubling policy, described in Sec. IV.1.
Figure 17: Average waiting time (left) and average age of the youngest end-to-end link (right) for a repeater chain with $n=9$ nodes (eight elementary links), according to the FN, SN, and random swap-asap policies, compared to the SN doubling policy. For small values of $p_{\ell}$, both the FN and SN swap-asap policies outperform the doubling policy in terms of waiting time. The SN swap-asap policy outperforms doubling for the average age for all values of $p_{\ell}$ up to $p_{\ell}=0.5$.
Our FN and SN swap-asap policies can outperform a doubling based multiplexing policy. This is what we show in Fig. 17. In terms of the average waiting time, the FN swap-asap policy outperforms the SN doubling policy for all values of link success probability $p_{\ell}$ up to $p_{\ell}=0.5$, including very low values of $p_{\ell}$, which is often most relevant in practice. It is important to mention here that this result is not true for all parameter regimes; for high values of $p_{\ell}$ and $m^{\star}$, the doubling policy can actually outperform our swap-asap policies. Our results are consistent with the results in previous works. Specifically, our results generalize those in Ref. [56] to the case of finite coherence times.
At the same time, our results also show the importance of ranking the links in an appropriate way, reflected in the fact that the random swap-asap policy and the SN swap-asap policies actually perform worse than the doubling policy in terms of average waiting time. We see very similar results in terms of the average age of the youngest end-to-end link. With respect to this figure of merit, the SN swap-asap policy outperforms all other policies discussed here, including the SN doubling policy.
Appendix B Noise model and decoherence
@@
For a proof, we refer to Ref. [45, Appendix E]. This implies that the fidelity of the state after $m$ time steps is $\frac{1}{4}(1+\mathrm{e}^{-2m/m_{1}^{\star}}+2\mathrm{e}^{-2m/m_{2}^{\star}})$.
@@
The state in (20) is then
$\sigma(m)$ $\coloneqq(\mathcal{N}_{2m^{\star},2m^{\star}}\otimes\mathcal{N}_{2m^{\star},2m^{\star}})^{\circ m}(\Phi^{+})$ (22)
$=\frac{1}{4}(1+3\mathrm{e}^{-\frac{m}{m^{\star}}})\Phi^{+}+\frac{1}{4}(1-\mathrm{e}^{-\frac{m}{m^{\star}}})(\Phi^{-}+\Psi^{+}+\Psi^{-})$ (23)
$=f(m)\Phi^{+}+\frac{1-f(m)}{3}(\Phi^{-}+\Psi^{+}+\Psi^{-}),$ (24)
where the fidelity $f(m)$ after $m$ time steps is equal to
$f(m)\coloneqq\frac{1}{4}(1+3\mathrm{e}^{-\frac{m}{m^{\star}}}).$ (25)
Note that we have chosen the value of $m_{2}^{\star}$ in Eq. (21) such that, at the cutoff time $m^{\star}$, the fidelity of the entangled qubit pair is $f(m^{\star})=\frac{1}{4}(1+3/e)\approx 0.5259$. We emphasize that our results can be applied to other choices of $m_{1}^{\star}$ and $m_{2}^{\star}$—in particular, choices that could take dephasing as the dominant source of noise [106], in order to make connections to prior works [61, 21, 24, 43, 31].
@@
$\rho_{(a,b,c)}\coloneqq(a+b)\Phi^{+}+(a-b)\Phi^{-}+c\Psi^{+}+c\Psi^{-},$ (27)
with $a=\frac{1}{4}(1+\mathrm{e}^{-2m/m_{1}^{\star}})$, $b=\frac{1}{2}\mathrm{e}^{-2m/m_{2}^{\star}}$, and $c=\frac{1}{4}(1-\mathrm{e}^{-2m/m_{1}^{\star}})$. As such, its entanglement can be characterized entirely by the quantity $a+b$, namely, the fidelity of the state with respect to $\Phi^{+}$. In particular, the state is entangled if and only if this fidelity is strictly greater than $\frac{1}{2}$; see, e.g., [55, Chapter 2]. Therefore, under the assumptions in Eq. (21), such that the fidelity is given by Eq. (25), we find that the state in Eq. (20) is entangled for all time steps $m\in\{0,1,2,\dotsc,m^{\star}\}$. Furthermore, inserting $F=\frac{1}{2}$ in Eq. (26), we find that the age at which entanglement is lost is $\lceil m^{\star}\log(3)\rceil$.
Appendix C Entanglement distillation with the BBPSSW protocol
@@
Note that these formulas apply only when the input states to the distillation protocol are so-called isotropic states, given by $f\Phi^{+}+(\frac{1-f}{3})(\Phi^{-}+\Psi^{+}+\Psi^{-})$, with $f\in[0,1]$ being the fidelity. Also, distillation should be considered useful only when it can improve upon the fidelity of the best of the two links. In other words, in order for distillation to be useful, we should have $F_{\text{distill}}(f_{1},f_{2})>\max\{f_{1},f_{2}\}$. In Fig. 18, we plot the $(f_{1},f_{2})$ region defined by this inequality.
From Ref. [79], we have the following relations, telling how the Bell states $\Phi^{\pm}$ and $\Psi^{\pm}$ are transformed by the BBPSSW protocol.
$\Phi^{+}$ $\Phi^{-}$ $\Psi^{+}$ $\Psi^{-}$
$\Phi^{+}$ $\Phi^{+}$ $\Phi^{-}$ 00 00
$\Phi^{-}$ $\Phi^{-}$ $\Phi^{+}$ 00 00
$\Psi^{+}$ 00 00 $\Psi^{+}$ $\Psi^{-}$
$\Psi^{-}$ 00 00 $\Psi^{-}$ $\Psi^{+}$
Here, the rows correspond to the first link, and the columns correspond to the second link. From this, it follows that the unnormalized state corresponding to success of the BBPSSW protocol, when both input states are of the form (24), is
$\sigma(m_{1})\otimes\sigma(m_{2})$ $\mapsto\left(f(m_{1})f(m_{2})+\frac{1-f(m_{1})}{3}\frac{1-f(m_{2})}{3}\right)\Phi^{+}+\left(f(m_{1})\frac{1-f(m_{2})}{3}+\frac{1-f(m_{2})}{3}f(m_{1})\right)\Phi^{-}$
$\qquad\qquad\qquad\qquad+2\frac{1-f(m_{1})}{3}\frac{1-f(m_{2})}{3}(\Psi^{+}+\Psi^{-})$ (31)
$\mapsto\left(\frac{10}{9}f(m_{1})f(m_{2})-\frac{1}{9}(f(m_{1})+f(m_{2}))+\frac{1}{9}\right)\Phi^{+}$
$\qquad\qquad\qquad\qquad+\left(-\frac{10}{9}f(m_{1})f(m_{2})+\frac{1}{9}(f(m_{1})+f(m_{2}))+\frac{8}{9}\right)\frac{1}{3}(\Phi^{-}+\Psi^{+}+\Psi^{-}),$ (32)
where in the last line we applied the isotropic twirling map $X\mapsto\mathrm{Tr}[\Phi^{+}X]\Phi^{+}+\frac{1-\mathrm{Tr}[\Phi^{+}X]}{3}(\Phi^{-}+\Psi^{+}+\Psi^{-})$ [107, Example 7.25], which is required in order to use the BBPSSW protocol recursively. Therefore, after normalization, the state is
$F_{\text{distill}}(f(m_{1}),f(m_{2}))\Phi^{+}+\frac{1-F_{\text{distill}}(f(m_{1}),f(m_{2}))}{3}(\Phi^{-}+\Psi^{+}+\Psi^{-}).$ (33)
We observe that this is a state of the form $\sigma(m^{\prime})$ in (24), for some age $m^{\prime}$. We can determine the age $m^{\prime}$ corresponding to the state after distillation via $m^{\prime}=f^{-1}(F_{\text{distill}}(f(m_{1}),f(m_{2})))$, with the expression for $f^{-1}$ given in Eq. (26). In particular, the updated age $m^{\prime}$ is given by:
@@
  • Pumping policy: Referred to as “scheme C” in Ref. [17], this policy proceeds as follows. Two links of fidelity $f_{0}$ are distilled, then a fresh link of fidelity $f_{0}$ is used to distill the previously distilled link, and so on. This policy is tailored to a practical situation in which fresh links are used to distill the surviving links of previous time steps, and thus the links being distilled do not necessarily have the same fidelity, as in the banded policy. If the initial fidelity satisfies $f_{0}>\frac{1}{2}$, the fidelity $f_{r}$ after $r\in\{1,2,\dotsc\}$ rounds of the pumping policy is given by the recurrence relation $f_{r}=F_{\text{distill}}(f_{r-1},f_{0})$. Below, in Theorem 1, we provide a closed-form, analytical solution to this recurrence relation.
C.1 The pumping policy
@@
We note that the limiting fidelity on the right-hand side of (55) is strictly less than one whenever $f_{0}<1$. This proves that, in contrast to the banded policy, with the pumping policy it is not possible to distill links to arbitrarily high fidelity, even with an infinite number of rounds.
C.2 Banding vs. pumping
@@
Figure 19: Comparison of the swap-distill (left) and distill-swap (right) policies for a three-node chain with $n_{ch}$ links connecting the neighboring nodes, as described in Sec. C.2. Here we use the banded policy to perform distillation. The swap-distill policy is not useful beyond initial age $m^{\star}/2$, whereas distill-swap can extend this threshold to $m^{\star}$.
@@
Figure 20: Comparison of the swap-distill (left) and distill-swap (right) policies for a three-node chain with $n_{ch}$ links connecting the neighboring nodes, as described in Sec. C.2. Now, instead of the banding policy, we use the pumping policy for distillation. As with banded purification, swap-distill is not useful beyond an initial age of $m^{\star}/2$, whereas distill-swap can extend this threshold, but now not quite up to $\approx m^{\star}$. This is a consequence of the fact that the pumping policy has a strict limit on the fidelity that can be achieved with an increasing number of channels, as shown in Eq. (55).
@@
Appendix D Distill-swap vs. swap-distill policy: analytical results
Here we analytically compare the performance of the distill-swap and swap-distill policies in the case of three nodes. To simplify the analysis, we restrict ourselves to the cases with two and four channels, and consider all elementary links to be active simultaneously with the same fidelity. The general result for an arbitrary number of channels is difficult to calculate analytically given the large number of possible trajectories that can lead to the successful generation of at least one end-to-end link. Seeing in Appendix C.2 that the banded policy performs better than pumping, in this section we look exclusively at the banded policy. The performance of the policies is evaluated by the success probability of establishing at least one end-to-end entangled link and by the average fidelity of an end-to-end link.
D.1 Two-channel case
Let $f_{0}$ be the initial fidelity of both the active links on each side given in terms of the initial age $m_{0}$ according to Eq. (25) as $f_{0}=f(m_{0})=\frac{1}{4}(1+3\mathrm{e}^{-m_{0}/m^{\star}})$.
Let us first consider the distill-swap policy. The probability of distillation being successful simultaneously on both sides followed by a successful entanglement swapping of the distilled links is given by
@@ D.1 Two-channel case
D.2 Four-channel case
@@
Figure 21: Comparison of the distill-swap and swap-distill policies for a three-node chain with four channels connecting the neighboring nodes, as described in Sec. D. We look at the probability of successfully producing at least one end-to-end entangled link (left) and the average age of that link (when successfully produced) (right). For distillation, we use the banded policy. We see again that the swap-distill policy is not useful beyond and initial age of $m^{\star}/2$, whereas distill-swap can extend this threshold up to $m^{\star}$ by using an increasing number of channels. $p_{sw}=50\%$ for these plots.
In Fig. 21, we show the results for four channels. It is clear from Fig. 21 that swap-distill has a higher probability of success compared to distill-swap when the initial age of the links is less than $m^{\star}/2$, beyond which it fails to succeed. Unlike distillation, success probability of entanglement swapping is independent of the quality of the used links. This is the primary reason for an asymmetry in the success rates of swap-distill and distill-swap policies. The probability of at least one entanglement swap being successful increases exponentially with the number of attempts available, whereas the probability of all distillations being successful falls with an increasing number of attempts, irrespective of the ages of the links involved. Thus, it is clear that if the aim is to generate at least one end-to-end link as quickly as possible, i.e., if the aim is to minimize waiting time, the swap-distill is the policy of choice. At the same time, distill-swap succeeds until a larger initial age threshold; thus, if the elementary links are of low fidelity to begin with, distilling first might be judicious. Also, in terms of fidelity enhancement (age reduction), distill-swap performs better. These conclusions are also consistent with the results in Fig. 19 and 20.
Appendix E Other policies for multiplexing based on swap-asap